OpenStudy (anonymous):

A sprinter in a race accelerates from rest at 1 m/s^2 for 10 s. She then continues east at her final speed for 45 s. How far does the sprinter run?

OpenStudy (anonymous):

Distance traveled is given as the product of speed and time. Acceleration describes how much speed increases each unit of time. In this case, 1 m/s^2 means the velocity of the sprinter increases 1 m/s every second. This happens for 10 seconds. So at the end of 10 seconds, the sprinter is now traveling (1) * (10) = 10 m/s To find the distance traveled during acceleration: Vf^2 = Vi^2 + 2a(Xf - Xi) 10^2 = 0 + 2(1)(Xf - 0) 100 = 2 Xf Xf = 50 m Now the sprinter continues for 45 seconds at this pace. x = (10) * (45) = 450 m So combined, the sprinter goes 50+450 = 500 m